### v3

@@ 6 Standardization Process
As wireless communication and AI integration is a focus of the industry’s research direction, important standardization organizations and industry alliances at home and abroad have launched research on integrating mobile communication networks and AI since 2017. These include the 3GPP, ITU, european telecommunication standardization institute (ETSI), china communication standardization association, IMT-2020 (5G) promotion group, IMT-2030 (6G) promotion group, etc. Currently, 3GPP has three technical specification groups (TSG): the CN, RAN, and service and system aspects (SA). Each TSG has 4 to 6 WG. The 3GPP SA WG2 (responsible for developing the overall 3GPP system architecture and services, including user equipment, access network, CT and IP multimedia subsystem system architecture and services) and the 3GPP RAN WG3 (responsible for the overall radio network architecture and the specification of protocols for the related network interfaces) already accomplished some specification work on data collection, data analytic, and associated procedures at CN and radio network. 3GPP SA WG5 has progressed in standardizing network management data analytics services for managing AI/ML Models and functions. ETSI’s zero-touch network and service management working group focuses on the research and analysis of technologies related to automated closed-loop operations, aiming to achieve a high degree of automation in network and service management. ETSI’s experiential networked intelligence (ENI) working group is focused on defining an experiential and aware network management architecture that uses AI to improve operators’ experience in network deployment and operations. Its goal is to use AI technology to improve the efficiency and automation of network management by defining a perception-adaptive-decision-execute control model. Figure 26 displays the application process of the AI model.
@@
In wireless access networks, 3GPP has identified typical application scenarios for SON and has applied ML techniques, successfully verifying its usefulness in optimizing network performance. In 2018, 3GPP SA5 Group set up a research project, “a study on SON for 5G,” aimed at the integration of SON and 5G network technology, studying to promote the integration of SON and 5G network infrastructure by strengthening the collection and application of wireless big data.
@@
The RAN1 Working Group identified the study item “Study on AI/ML for NR air interface” in the Rel-18 phase [256]. The SI explores how AI/ML algorithms can enhance air interface performance (throughput, robustness, accuracy, or reliability) and reduce overhead. The main application scenarios of this research include AI-based CSI feedback, AI-based beam management, AI-based positioning enhancement, and life cycle management of AI/ML model/functionality for the air interface.
The 3GPP RAN2 Working Group identified the “Study on AI/ML for mobility in NR” research project in December 2023 [257]. The study project aims to evaluate potential benefits and potential specification impacts of AI/ML-aided mobility for network-triggered L3-based hand off, considering the following aspects: AI/ML-based RRM measurement and event prediction for L3 Mobility, HO failure/RLF prediction, and Measurement events prediction.
@@
The 3GPP RAN3 Working Group identified the study item “Study on further enhancement for data collection” in July 2020 [258]; the SI aims to enable an intelligent functional architecture for RAN through use cases to enhance data collection further and identify potential standardization implications for current NG-RAN nodes and interfaces. In February 2022, the SI was completed, and the intelligent typical functional architecture on the RAN side, input, output, feedback, and potential standard impact of three high-priority use cases (e.g., network energy saving, load balancing, and mobility optimization) were written into TR 37.817. The intelligent universal functional architecture on the RAN side includes functions such as data collection, model training, model inference, and action execution. The 3GPP RAN3 Working Group identified the “AI/ML for NG-RAN” work item [259]. The WI aims to incorporate data collection enhancements and signaling support specified in existing NG-RAN interfaces and architectures to enable AI/ML-based network energy saving, load balancing, and mobility optimization.
@@
The TSG SA (responsible for the overall architecture and service capabilities of 3GPP systems and the cross-3GPP TSG coordination) identified the Study item, “3GPP AI/ML Consistency Alignment,” in June 2024 [260]. The SI will investigate ongoing AI/ML work in TSG CT, TSG RAN and TSG SA working groups and identify instances of any potential misalignment and/or inconsistencies. TSG SA leads the study in close collaboration and inputs from TSG CT and TSG RAN. The technical report of this SI will be finalized by June 2025.
@@
In addition to the 3GPP-defined NWDAF network element, the 5G-Monarch project in Europe introduced RAN-DAF, an independent AI analysis function explicitly designed for control unit plane in 5G NR, for data analysis and decision-making. As a network element for AI and data analysis on the RAN side, RAN-DAF can monitor and collect UE and RAN data, and AI can prioritize local processing based on these real-time data to solve the need for rapid response of operations such as wireless resource management. MoNArch recommends that the RAN-DAF pass information to the controller on the RAN side to collaborate with the RCA to optimize wireless side performance, such as slice-sensitive selection, live wireless resource control, cross-slice resource management, and so on.
@@
In 2017, ETSI established the ENI Working Group dedicated to applying AI technologies to network operations. Improve the experience of operator network deployment and operations with a closed-loop AI mechanism based on context-aware and data-driven strategies. Since its inception, ENI has published several versions of specifications and reports, including system architectures with context-aware policy management, data processing mechanisms, and hierarchical evaluation methods for network AI. In 2019, ETSI ENI published the first edition of this research report and standard and launched the second phase of the closed-loop control study for real-time networks.
In recent years, standardization organizations have also focused on the research and preparation of 6G. In June 2019, the China IMT-2030 (6G) promotion group was established to gather the strength of Chinese industry, universities and research institutes, promoting China’s 6G mobile communication technology research and carrying out international exchanges and cooperation and successively releasing several white papers and research reports such as the data format and model proposal of mobile communication and AI integration. In June 2023, the 44th meeting of the ITU-R working group 5D was held in Geneva, Switzerland, and the recommendation on the framework and overall objectives of the IMT for 2030 and Beyond was completed. This proposal aims to set a framework and overall goals for the development of IMT in 2030 and beyond, involving a global 6G vision consensus, 6G goals and trends, typical 6G scenarios and capability indicators system, including the convergence of AI and communication scenarios, and define AI-related indicators. On September 12, 2024, at the 105th Plenary meeting held in Melbourne, Australia, 3GPP officially launched the first 6G Study Item - “6G use cases and Service Requirements”. This initiative formally marks the 6G technology from the pre-research stage into the substantive standardization stage.
@@
The vigorous development of AI standardization will significantly promote the integration of communication and AI, as shown in Figure 27. AI still has a way to go from standardization to landing, which is also the main research direction of many standardization organizations in the future.
@@ 7 Challenges of AI and communication for 6G
Integrating AI into 6G networks presents a dual-edged sword, offering significant advantages and notable challenges. On the one hand, AI enhances network efficiency, enabling intelligent resource allocation, real-time data analysis, and adaptive communication strategies. Semantic communication, for instance, leverages AI to interpret and convey the intended meaning of messages, reducing bandwidth usage and improving transmission accuracy. Generative AI further contributes by creating predictive models that enhance network performance and user experience.
On the other hand, these advancements introduce complexities. The challenges of AI for the network include the strain on bandwidth and storage due to complex models. Conversely, the challenges of the NET4AI require robust infrastructure to support advanced functionalities. The reliance on AI raises concerns about sustainability, as increased energy demands impact the environmental footprint. The requirement for real-time AI necessitates ultra-low latency, demanding sophisticated computational infrastructure. Security and privacy are at risk from potential attacks on sensitive data. Lastly, human-AI interactive collaboration presents ethical challenges, necessitating transparent oversight to ensure trust and accountability. This section focuses on the multifaceted challenges faced by AI and communication for 6G from the abovementioned perspectives.
@@ 7.1 Challenges of AI4NET
The reliability of AI in networks poses a significant challenge due to the dynamic, heterogeneous, and often unpredictable nature of 6G communication systems [261]. AI models, which rely heavily on data-driven learning, can struggle to maintain consistent performance when exposed to conditions that deviate from their training environments. For instance, unexpected network anomalies, rare edge cases, or sudden shifts in user behavior patterns can lead to errors or even system failures [262]. Moreover, the non-deterministic nature of many AI algorithms makes it difficult to guarantee predictable outcomes, especially in time-sensitive or mission-critical applications. Factors such as data sparsity, incomplete information, and the presence of adversarial inputs further undermine the reliability of AI, as these systems may produce incorrect predictions or decisions under such conditions. Additionally, the distributed and decentralized nature of 6G networks introduces further complexity, as AI models must operate across diverse nodes with varying resource constraints and environmental conditions. This interconnectedness increases the risk of cascading failures, where an error in one part of the system can propagate and disrupt the entire network. These challenges highlight the fundamental difficulty in ensuring that AI systems remain reliable, robust, and consistent in the face of evolving and unpredictable network demands.
@@
The stability of AI in networks is a critical challenge, as it directly impacts the consistency and predictability of AI-driven decisions in dynamic and complex environments [263]. Stability refers to an AI system’s ability to produce reliable outputs when exposed to variations in input data, environmental conditions, or system parameters. In networked systems, this becomes particularly challenging due to the heterogeneity and unpredictability of real-world conditions [264]. For instance, fluctuating network traffic, sudden surges in user demand, or unexpected hardware failures can introduce variations that destabilize AI models. This instability can manifest as erratic predictions, oscillating decision patterns, or even complete system breakdowns. Furthermore, the non-linear and highly sensitive nature of many AI algorithms, such as DL models, makes them prone to overreacting to small changes in input data, leading to cascading errors. In distributed or decentralized networks, such instability can propagate across nodes, amplifying its impact and potentially disrupting the entire system. Additionally, the lack of transparency in many AI models (e.g., black-box models) complicates the diagnosis and mitigation of instability, as it is often unclear how or why a model’s output changes under varying conditions. These challenges highlight the inherent difficulty in ensuring stable and predictable AI performance in the face of the complex, evolving, and often adversarial nature of modern networks.
@@
The generalization problem is widely prevalent in integrating AI with communication systems [265, 266, 267, 268]. Network models with poor generalization capabilities cannot often extend their learned knowledge to other variable domains. Consequently, such models favor one or a few data types, resulting in suboptimal performance when reasoning on different data types. This issue becomes particularly critical when the model is exposed to diverse datasets on which it was not explicitly trained, leading to a significant decline in its effectiveness. The application of AI in wireless communications imposes even higher demands on model generalization [269, 270, 271, 272, 273]. In wireless environments, users are situated in varying contexts, with factors such as the relative position to BSs and user mobility playing significant roles. This variability means that pre-trained AI models may lose effectiveness when confronted with new data that deviate from their training sets. In the dynamic and complex wireless environments characteristic of 6G networks, models with poor generalization capabilities will lack the potential for deployment and widespread use. Therefore, enhancing the generalization ability of AI models within 6G networks is a critical issue requiring industry and academia dedicated attention. Addressing this challenge will ensure that AI models maintain high performance and adaptability across diverse and evolving wireless scenarios, thus facilitating their successful integration into future communication systems.
@@
DL is essentially a black box due to its highly nonlinear nature. The model learns through neurons, convolution kernel and other ways. People cannot understand the meaning, importance and fluctuation range of parameters in the neural network through basic statistical assumptions like linear regression parameters, which leads to unexplainability. For example, we only focus on the encoder and decoder output if we use AI for CSI feedback. We do not have a concrete understanding of how AI performs feature extraction and recovery, leading to the randomness of model training and inference performance that is difficult to control and increased unreliability against harsh, extreme environments. Interpretable AI is not only a technical issue but also a complex issue involving trust, fairness, regulation, and human relations. Improving the explainability of AI and helping human users understand the decision-making process, working principle and potential bias of AI to build trust in AI is necessary to promote the broad application of AI in communication and even in various fields. The millions or even billions of parameters in DNNs and their nonlinear activation function form a highly complex decision boundary. This complexity makes the model’s decision-making process difficult to understand intuitively, thus posing significant challenges for interpretability.
Robustness of AI models
Robustness, derived from the concept of cybernetics, plays a central role in AI as it measures whether an AI model can maintain its efficacy and reliability in the face of real-world complexity and uncertainty. A robust AI model can keep the stability of its output in the face of minor deviations from the data. This property is critical in determining whether the model suits real-world scenarios. In modern wireless communication systems, AI technologies such as DL are playing an increasingly important role[274]. For example, in the field of semantic communication, DNNs are widely used for critical tasks such as semantic coding, decoding and understanding[275]. However, the fragility of DL models and the open nature of wireless channels make these AI-powered communication systems extremely vulnerable to physical layer adversarial attacks. Attackers can mislead the semantic interpretation on the receiving end by injecting carefully designed perturbation signals, causing catastrophic consequences in certain critical domains that rely on highly reliable communication (e.g., autonomous driving, smart healthcare, etc.). In this context, the robustness of AI algorithms in communication systems is particularly important. Ensuring the robustness of AI algorithms in communication systems has become a critical issue.
@@ 7.2 Challenges of NET4AI
Constructing an AI execution environment in wireless networks inherently entails significant dynamic variations. The interference and loss in the wireless connection between terminals and BSs are impacted by the surrounding urban environment and changes in the operational states of other BSs, subsequently affecting the uplink and downlink connection rates. At the same time, terminals exhibit high mobility, potentially moving from the center to the edge of a cell, hand off to another cell, or even leaving the wireless coverage area. Additionally, distinct from dedicated computing resources in IT clouds, wireless networks carry conventional mobile connectivity tasks, with the mobile traffic load on individual network elements continually fluctuating. BSs alternate between busy and idle periods, and terminals enter an idle state when there are no uplink or downlink services. These variations contribute to the dynamic nature of mobile wireless networks, resulting in slow model training convergence and adversely affecting model inference performance. It can lead to the failure or loss of partial information, such as gradients or losses, during AI model computations, further slowing down convergence and, in severe cases, causing divergence and failure to converge. Acknowledging the influence of these dynamic wireless transmission factors, the authors of [276] designed a mitigation strategy for unreliable communication nodes. This approach leverages dynamic wireless channel state predictions to estimate the training completion latency for each user equipment’s local model. It avoids unreliable wireless communication nodes, prioritizing nodes with stable and high-speed wireless connections for training participation. When unstable or disconnected device nodes occur during training, the strategy advocates for their exclusion to ensure continuous and efficient AI training. Simulation results demonstrate that avoiding unreliable nodes facilitates stable convergence during AI model training.
@@
In wireless communication networks, a myriad of heterogeneities exist, encompassing communication heterogeneity arising from the varying geographical locations of nodes and the time-varying nature of communication network links; computational heterogeneity due to differences in the number, manufacturing processes, and architectures of node chips; data heterogeneity in terms of the volume and distribution of data possessed by nodes, influenced by factors such as storage capacity and application scenarios; and model heterogeneity stemming from varying task objectives that lead to the training of diverse models. These heterogeneities pose challenges to achieving efficient distributed learning. This diverse heterogeneity influences both AI performance and convergence. Data heterogeneity results in disparities among local gradients computed by AI models, causing inconsistency in the direction of aggregated gradients, which can lead to model divergence or even non-convergence. Differences in computational power and communication capabilities lead to variations in AI model computation and information transmission efficiency, giving rise to the straggler effect. Furthermore, in AI learning, when the number of user devices is sufficiently large or the local models are adequately complex, the consumption of wireless resources during transmission becomes significant.
Discreteness of Networks
The discrete nature of 6G mobile networks, characterized by highly heterogeneous and dynamically changing environments, poses significant challenges to the application of AI. This discreteness manifests through the rapid and unpredictable variations in network topology, user mobility patterns, and service demands, leading to inconsistent and fragmented data streams. Such conditions impede the ability of AI models to maintain accuracy and reliability, as the models must continuously adapt to a multitude of scenarios with limited contextual information. The resulting variability can degrade the performance of AI algorithms, particularly those relying on stable and predictable data inputs, thereby affecting tasks such as predictive analytics, resource allocation, and anomaly detection. These challenges necessitate a concerted effort from academia and industry to develop AI models with enhanced robustness and generalization capabilities. This involves advancing research in adaptive learning techniques, transfer learning, and FL for the academic community to handle better the sporadic and diverse data inherent in 6G networks. For the industry, there is a pressing need to design and implement resilient AI-driven network management systems capable of real-time adaptation to fluctuating conditions. Addressing these challenges satisfactorily will be crucial for leveraging the full potential of AI in optimizing the performance and efficiency of 6G mobile networks.
@@
The large scale and complexity of 6G networks significantly affect the deployment and performance of AI systems. The many connected devices, diverse applications, and heterogeneous network environments introduce unprecedented data volume, variety, and velocity levels. AI models must process and analyze this extensive data in real time to make effective decisions, which is a daunting task due to the high-dimensional and dynamic nature of the information. The large-scale deployment of AI in 6G networks also necessitates distributed and decentralized processing architectures, complicating data aggregation and synchronization. Furthermore, the intricate interplay between various network components, such as edge and CNs, requires AI models to handle multifaceted interactions and dependencies, leading to increased computational overhead and latency. Consequently, AI algorithms’ scalability and efficiency are tested, making it challenging to maintain high performance and reliability across the entire network. Addressing these challenges demands innovative approaches from both academia and industry. From the perspective of academia, there is a need to advance research in scalable AI techniques, including distributed learning, edge computing, and hierarchical models, to efficiently handle the massive and complex data in 6G networks. Developing AI algorithms capable of real-time processing and decision-making with minimal latency is crucial. Additionally, robust frameworks for integrating AI with existing network infrastructures must be explored to ensure seamless operation. The industry should focus on implementing scalable AI-driven solutions that can dynamically adapt to the network’s complexity and scale. This involves designing intelligent network management systems that can autonomously optimize resources, enhance security, and improve user experiences. Collaborative efforts between academia and industry will be essential to develop and deploy AI technologies that can fully leverage the capabilities of 6G networks, ensuring robust performance and efficient operation in an increasingly complex and large-scale environment.
@@
In large-scale and dynamic edge federated networks, the localized model training and transmission may not always succeed due to constraints on mobile devices’ power and computational resources, bandwidth limitations, wireless channel impairments, etc., directly compromising the performance of FL regarding model accuracy and convergence speed. Consequently, evaluating the trade-off relationship between cognitive performance and multi-dimensional resource allocation becomes imperative upon introducing distributed learning. The authors of [277] focus on FL-enabled edge intelligence networks by leveraging Poisson Point Processes to construct an analytical model. The work delves into three specific scenarios, each deriving mathematical relationships between FL performance, consumed computational resources, and communication resources by analyzing the impact of Signal-to-Interference-plus-Noise Ratio, SNR, and faulty transmissions on local and global models. This provides a theoretical foundation and basis for optimizing the framework design of intelligent access networks and allocating communication, computing power, and data resources.

### v4

@@ 6 Standardization Process
As wireless communication and AI integration is a focus of the industry’s research direction, important standardization organizations and industry alliances at home and abroad have launched research on integrating mobile communication networks and AI since 2017. These include the 3GPP, ITU, European Telecommunication Standardization Institute (ETSI), China Communication Standardization Association, IMT-2020 (5G) promotion group, IMT-2030 (6G) promotion group, etc. Currently, 3GPP has three technical specification groups (TSG): the CN, RAN, and service and system aspects (SA). Each TSG has 4 to 6 WG. The 3GPP SA WG2 (responsible for developing the overall 3GPP system architecture and services, including user equipment, access network, CN and terminals (CT), and internet protocol (IP) multimedia subsystem system architecture and services) and the 3GPP RAN WG3 (responsible for the overall radio network architecture and the specification of protocols for the related network interfaces) already accomplished some specification work on data collection, data analytic, and associated procedures at CN and radio network. 3GPP SA WG5 has progressed in standardizing network management data analytics services for managing AI/ML Models and functions. ETSI’s zero-touch network and service management WG focuses on the research and analysis of technologies related to automated closed-loop operations, aiming to achieve a high degree of automation in network and service management. ETSI’s experiential networked intelligence (ENI) WG is focused on defining an experiential and aware network management architecture that uses AI to improve operators’ experience in network deployment and operations. Its goal is to use AI technology to improve the efficiency and automation of network management by defining a perception-adaptive-decision-execute control model. Figure 23 displays the application process of the AI model.
6.1 3GPP
@@
3GPP has identified typical application scenarios for self-organizing networks (SON) and has applied ML techniques, successfully verifying its usefulness in optimizing network performance. In 2018, 3GPP SA5 Group set up a research project, “A study on SON for 5G”, aimed at the integration of SON and 5G network technology, studying to promote the integration of SON and 5G network infrastructure by strengthening the collection and application of wireless big data.
@@
The RAN1 WG initiated the study item “Study on AI/ML for NR air interface” as part of the Release-18 phase [282]. The study item explores how AI/ML algorithms can enhance air interface performance (throughput, robustness, accuracy, or reliability) and reduce overhead. The main application scenarios of this research include AI-based CSI feedback, AI-based beam management, AI-based positioning enhancement, and life cycle management of AI/ML model/functionality for the air interface. Similarly, the 3GPP RAN2 WG launched the “Study on AI/ML for mobility in NR” research project in December 2023 [283]. The study project evaluates the potential benefits and specification impacts of AI/ML-aided mobility for network-triggered L3-based handoff. The project addresses several critical aspects, including AI/ML-based radio resource management measurement and event prediction for L3 Mobility, handoff/radio link failure prediction, and Measurement events prediction.
@@
The 3GPP RAN3 WG identified the study item “Study on further enhancement for data collection” in July 2020 [284]. This study item aimed to enable an intelligent functional architecture for RAN. By leveraging use cases, it targeted to enhance data collection and identify potential standardization implications for NG-RAN nodes and interfaces. In February 2022, the study item was completed, and the intelligent typical functional architecture on the RAN side, input, output, and feedback. The potential standard impact of three high-priority use cases (e.g., network energy saving, load balancing, and mobility optimization) was incorporated into TR 37.817. The intelligent universal functional architecture on the RAN side includes functions, such as data collection, model training, model inference, and action execution. Recently, the 3GPP RAN3 WG has launched the “AI/ML for NG-RAN” work item [285]. The WI aims to incorporate data collection enhancements and signaling support specified in existing NG-RAN interfaces and architectures to enable AI/ML-based network energy saving, load balancing, and mobility optimization.
@@
The TSG SA (responsible for the overall architecture and service capabilities of 3GPP systems and the cross-3GPP TSG coordination) launched the study item, “3GPP AI/ML Consistency Alignment,” in June 2024 [286]. The study item investigates ongoing AI/ML work in TSG CT, TSG RAN, and TSG SA WGs, and identifies the instances of any potential misalignment and/or inconsistencies. TSG SA leads the close collaboration and takes inputs from TSG CT and TSG RAN. The technical report of this study item will be finalized by June 2025.
6.2 5G-MoNAch
@@
The 5G-mobile network architecture (5G-MoNArch) project in Europe introduced RAN-DAF, an independent AI analysis function explicitly designed for control unit plane in 5G NR, for data analysis and decision-making [287]. As a network element for AI and data analysis on the RAN side, RAN-DAF can monitor and collect UE and RAN data, and AI can prioritize local processing based on these real-time data to solve the need for rapid response of operations such as wireless resource management. MoNArch recommends that the RAN-DAF pass information to the controller on the RAN side to collaborate with the RCA to optimize wireless side performance, such as slice-sensitive selection, live wireless resource control, cross-slice resource management, and so on.
6.3 ITU-T
@@
6.4 ETSI
In 2017, ETSI established the ENI WG dedicated to applying AI technologies to network operations. Improve the experience of operator network deployment and operations with a closed-loop AI mechanism based on context-aware and data-driven strategies. Since its inception, ENI has published several versions of specifications and reports, including system architectures with context-aware policy management, data processing mechanisms, and hierarchical evaluation methods for network AI. In 2019, ETSI ENI published the first edition of this research report and standard and launched the second phase of the closed-loop control study for real-time networks.
@@ 6.4 ETSI
6.5 Collaborative Efforts in 6G
In recent years, standardization organizations have also focused on the research and preparation of 6G. In June 2019, the China IMT-2030 (6G) promotion group was established to gather the strength of Chinese industry, universities, and research institutes, promoting China’s 6G mobile communication technology research and carrying out international exchanges and cooperation and successively releasing several white papers and research reports such as the data format and model proposal of mobile communication and AI integration. In June 2023, the 44th meeting of the ITU-R WG 5D was held in Geneva, Switzerland, and the recommendation on the framework and overall objectives of the IMT for 2030 and Beyond was completed. This proposal aims to set a framework and overall goals for the development of IMT in 2030 and beyond, involving a global 6G vision consensus, 6G goals and trends, typical 6G scenarios, and capability indicators system, including the convergence of AI and communication scenarios, and define AI-related indicators. On September 12, 2024, at the 105th Plenary meeting held in Melbourne, Australia, 3GPP officially launched the first 6G study item – “6G use cases and Service Requirements”. This initiative formally marks the 6G technology from the pre-research stage into the substantive standardization stage. The vigorous development of AI standardization will significantly promote the integration of communication and AI, as shown in Figure 24. AI still has a way to go from standardization to landing, which is also the main research direction of many standardization organizations in the future.
@@ 7 Challenges of AI and communication for 6G
This section delves into multiple challenges confronted by AI and communication in 6G. Herein, AI4NET and NET4AI represent different perspectives of AI empowering networks and networks supporting AI[288], respectively. Sustainability cannot be overlooked in the integration process of 6G and AI. The demands for real-time AI, security and privacy, and human-AI collaboration also bring challenges to the development of AI and communication in the 6G era.
@@ 7.1 Challenges of AI4NET
The reliability of AI in networks poses a significant challenge due to the dynamic, heterogeneous, and often unpredictable nature of 6G communication systems [289]. AI models, which rely heavily on data-driven learning, can struggle to maintain consistent performance when exposed to situations that deviate from their training environments. For instance, unexpected network anomalies, rare edge cases, or sudden shifts in user behavior patterns can lead to errors or even system failures [290]. Additionally, the distributed and decentralized nature of 6G networks introduces further complexity, as AI models must operate across diverse nodes with varying resource constraints and environmental conditions. This interconnectedness increases the risk of cascading failures, where an error in part of the system can propagate and disrupt the entire network. An approach based on model-driven engineering principles was proposed in [291], which allows ML experts to schedule the execution of drift-detecting algorithms on a computing cluster. The author of [292] proposed a distributed parallel modeling and monitoring framework for plant-wide processes with big data, where the multilevel monitoring indexes and fault contribution indexes are established based on the Bayesian fusion algorithm.
@@
Stability refers to an AI system’s ability to produce reliable outputs when exposed to variations in input data, environmental conditions, or system parameters. The stability of AI in networks is a critical challenge, as it directly impacts the consistency and predictability of AI-driven decisions in dynamic and complex environments [293]. This is particularly challenging in networked systems due to the heterogeneity and unpredictability of real-world conditions [294]. For instance, fluctuating network traffic, sudden surges in user demand, or unexpected hardware failures can introduce variations destabilizing AI models. The simulation results of [295] indicate that reducing the number of model parameters can enhance the model’s robustness against maximum adversarial attacks, while increasing the number of model parameters can improve the model’s robustness against minimum adversarial attacks. The authors of [296] studied the robustness of mainstream link prediction methods under various network attacks and discovered that the link prediction method with high performance probably has low attack robustness.
@@
The generalization problem is widely prevalent in integrating AI with communication systems [297]. The application of AI in wireless communications imposes even higher demands on model generalization [298]. In wireless environments, users are situated in varying contexts, with their relative position to the BSs and mobility playing significant roles. This variability means that pre-trained AI models may lose effectiveness when confronted with new data that deviate from their training sets. In the dynamic and complex wireless environments of 6G networks, models with poor generalization capabilities will lack the potential for deployment and widespread use. Data enhancement or meta-learning can potentially improve model generalization. The scenario-adaptive meta-learning for mmWave beam alignment method was proposed to improve the domain adaption capabilities of beam alignment problem in [299]. A fast adaptive channel prediction meta-learning technique combined with a denoising process was proposed in [300] to improve the generalization of the predictor.
@@
DL is often treated as a black box due to its highly nonlinear nature. Humans can hardly understand the meaning, importance, and fluctuation range of parameters in the neural network through basic statistical assumptions like linear regression parameters, leading to unexplainability. Improving the explainability of AI helps humans understand the decision-making process, working principle, and potential bias of AI, and build trustworthy AI [301, 302]. However, millions or even billions of parameters in DNNs and their nonlinear activation function form a highly complex decision boundary, hindering the interpretability of a model’s decision-making process. Through the gradient derivation and visualization of the model training process, the interpretability of the model can be effectively improved. The authors of [303] deduced the gradient expression of the meta-model to improve the interpretability in the process of backpropagation and promoted the application of meta-learning. t-SNE visualization was proposed in [304] to visualize the representations of data points. t-SNE is an embedded model that can map data from a high-dimensional space to a low-dimensional space and retain the local characteristics of datasets.
@@ 7.2 Challenges of NET4AI
Constructing an AI execution environment in wireless networks inherently entails significant dynamic variations. The interference and loss in the wireless connections between terminals and BSs are impacted by the surrounding environment, changes in the operational states of other BSs, and the mobility of terminals. These variations contribute to the dynamic nature of mobile wireless networks, resulting in slow model training convergence and adversely affecting model inference performance, and even causing model divergence. A mitigation strategy for unreliable communication nodes was designed in [158] to prevent unreliable nodes from contributing to stable convergence during the AI training process. In response to the transmission errors caused by the dynamics of networks, the authors of [305] modified the global aggregation method to counteract the model drift resulting from data packet errors. The theoretical relationship between dynamic unreliable communication channels and collaborative learning was also established in [306].
@@
There is a myriad of heterogeneities in wireless networks: communication heterogeneity arising from the geographical locations of nodes and the time-varying nature of communication links; computational heterogeneity due to differences in the number, manufacturing processes, and architectures of node chips; data heterogeneity in the volume and distribution of data possessed by nodes; and model heterogeneity stemming from varying task objectives. Data heterogeneity results in disparities among local gradients computed by AI models. Differences in computational power and communication capabilities lead to variations in model training and transmission efficiency, giving rise to the straggler effect. In [159], the sampling of training nodes in collaborative learning was optimized from the dimensions of communication, data, and computation, achieving efficient training on heterogeneous edge networks. For the heterogeneity of different models, the authors of [307] provided approximate rules for the decision boundaries of each model to bridge the gap caused by model heterogeneity, achieving more accurate model precision.
@@
The large scale and complexity of 6G networks significantly affect the deployment and performance of AI systems. In some cases, AI models must train and infer in real time to make effective decisions, which is a daunting task due to the high dimension and dynamic nature of the network data. The deployment of AI in 6G networks also necessitates decentralized architectures, complicating data aggregation and synchronization. Furthermore, the intricate interplay between various network components, such as edge and CN, requires AI models to handle multifaceted interactions and dependencies. To address these challenges, there is a need to advance research in scalable AI techniques. The authors of [308] proposed a hierarchical RL framework based on human prior knowledge. Through process decomposition, stage transformation, key feature selection, and a policy gradient with a parameter-based exploration method, it provides insights in solving complex operations. In response to the differentiated data distribution, a clustering collaborative learning framework was proposed in [309], where the server dynamically determines the optimal number of clusters by iteratively performing incremental clustering. Another focus could be on implementing scalable AI solutions that can dynamically adapt to the network complexity and scale. The authors of [310] used MARL to learn an adaptive fully distributed collaboration strategy for each collaborative node in complex wireless networks.
@@
The localized model training and transmission may not always succeed due to constraints on mobile devices’ power and computational resources, bandwidth limitations, and wireless channel impairments, directly compromising the performance of FL regarding model accuracy and convergence speed. Consequently, evaluating the trade-off relationship between cognitive performance and multi-dimensional resource allocation becomes imperative upon introducing distributed learning. For FL-enabled edge intelligence, the authors of [311] derived mathematical relationships between FL performance, consumed computational resources, and communication resources, and provided a theoretical foundation for the design of intelligent access networks. To meet the low-latency requirements of multimedia MEC services, a task offloading and resource allocation algorithm based on the double DQN was proposed in [312], which assigns appropriate task volumes to different devices and optimizes their communication and computing resource settings via BS. Moreover, the authors of [313] introduced a distributed resource allocation method. Leveraging meta-federated RL, each device can autonomously optimize its transmission power and channel usage according to wireless environments.